What are the dimensions or units of A in the equatice below that will make the equation unitically corect? Where has dimensions of LVIT1 or units of (myts
In which of the following cases does a car have a negative velocity and a positive acceleration? A car that is traveling in the O a.-x direction at a constant 20 m/s O b.-x direction decreasing in speed O c x direction increasing in speed O d. x direction increasing in speed o e x direction decreasing in speed
You drive 4 km at 30 km/hr and then another 4 km at 50 km/h. What is your average speed for the whole 8 km trip? O a less than 40 km/h O b. equal to 40 km/h O c. more than 40 km/h O d none of the above
Answer #1

(1) it is given that

a = At3

than -

A = a/t3 = m/s2×s3

A = m/s5

(2) option(b)

velocity is negative so velocity direction is -x direction. but acceleration is positive so speed is decreasing in -x direction.

(3) option(a) less than 40km/h

average velocity is (v) = 2v1v2/(v1+v2)

v = 2×30×50÷(50+30)

= 37.5km/h